# QMeCha: quantum Monte Carlo package for fermions in embedding environments

### v1

@@ Abstract
We present the first open access version of the QMeCha (Quantum MeCha /’m”ke/) code, a quantum Monte Carlo (QMC) package developed to study many-body interactions between different types of quantum particles, with a modular and easy-to-expand structure. The present code has been built to solve the Hamiltonian of a system that can include nuclei and fermions of different mass and charge, e.g. electrons and positrons, embedded in an environment of classical charges and quantum Drude oscillators. To approximate the ground state of this many-particle operator, the code features different wavefunctions. For the fermionic particles, beyond the traditional Slater determinant, QMeCha also includes Geminal functions such as the Pfaffian, and presents different types of explicit correlation terms in the Jastrow factors. The classical point charges and quantum Drude oscillators, described through different variational ansätze, are used to model a molecular environment capable of explicitly describing dispersion, polarization, and electrostatic effects experienced by the nuclear and fermionic subsystem. To integrate these wavefunctions, efficient variational Monte Carlo and diffusion Monte Carlo protocols have been developed, together with a robust wavefunction optimization procedure that features correlated sampling. In conclusion, QMeCha is a massively parallel code introduced here to explore quantum correlation effects in mixed systems with thousands of fermions and bosonic particles, beyond what was previously accessible to other reference methods.
@@ I Introduction
The most successful of these computational approaches fall into three main categories. These include wavefunction-based methods, such as Hartree-Fock (HF)[3, 4, 5, 6], Configuration Interaction[14, 15, 16] (CI) and Coupled Cluster[17, 18, 19, 20] (CC) multideterminantal expansions, complete active-space (CAS) theories[21]; the long list of perturbative approaches like Møller-Plesset (MP) perturbation theory[22] and symmetry-adapted perturbation theory (SAPT)[23]; and the density-based methods, such as Density Functional Theory (DFT)[24, 25].
@@
The long list of modern QMC methods can be divided into two main categories. The first category includes the methods that are based on the stochastic sampling of the system’s degrees of freedom, such as the particles’ coordinates, and includes the most famous variational Monte Carlo (VMC)[26, 27], diffusion Monte Carlo (DMC)[28] and Lattice-Regularized diffusion Monte Carlo (LRMDC)[31]. The second category, on the other hand, contains all the QMC methods based on the sampling of the configurational space, such as Auxiliary-Field QMC[32], FCI QMC[33, 34], and many others[28, 35].
Within this vast ensemble of methods and codes[36], in this work, we present the first version of the QMeCha (Quantum MeCha /’m”ke/)  [^1] The term mecha /’m”ke/ is the abbreviation of mechanical, which in Japanese manga and anime culture refers to all the piloted mechanical devices (robots, cars, airships, computers, and others) introduced in the science fiction genre. QMC package, initiated in 2017 as a flexible tool to study many-particle Hamiltonians with different wavefunction’ approximations. In the years that followed, the code progressively grew to describe interactions between fermions of different charge and mass, such as positrons, muons and anti-muons, interacting with molecules[38, 39], and to include a QMC embedding procedure in which a molecular system of nuclei and fermions can be immersed in an environment approximated by point charges and quantum Drude oscillators (QDOs)[40, 41, 42]. The full Hamiltonian that comprises the two types of quantum particles, ie fermions and drudons, is integrated over one correlated wavefunction that is constructed as the product of various fermionic terms, such as the Slater determinant, and Jastrow factors. Within the code, the observables are computed through an efficient implementation of the VMC method, and through an improved size-consistent DMC algorithm, that reduces finite time-step errors when computing energy differences.
@@
Finally, QMeCha has been used to tackle three main sets of applications applications reported in Section VI. The first application focuses on the description of van der Waals (vdW) interactions in macromolecules[43], where the DMC calculations are used as state-of-the-art references together with the LNO-CCSD(T)[44, 45, 46, 47] method, to construct a dataset for the description of pocket-ligand interactions in proteins. The second application is related to positronic chemistry, an emerging field dedicated to the investigation of the still vastly unexplored bound metastable states of positrons with atomic and molecular systems, before electron-positron pair annihilation[48, 49, 50]. It has been shown that QMC methods are a promising tool for studying electron-positron bound states, because they can explicitly describe the electrostatic correlation effects between electron–positron pairs through wavefunctions including the two particle distances[51, 52, 53, 38, 54, 55, 56], unlike many quantum-chemical post-HF methods (like MP, CC, and CI) that have to rely on extremely large single-particle basis sets. The third set of applications is that of the El-QDO embedding method[40, 41, 42] first introduced in ref. 40 to efficiently capture the quantum correlation effects between an environment modelled through QDOs and point charges[57, 58, 59] and a molecular subsystem of interest. In this application we have focused on the evaluation of solvation energies, excitation energies and bond energy variations of a set of dimers in water clusters.[40, 42]
@@ II.1 Fermionic Hamiltonian
QMeCha was first implemented to study non-periodic molecular systems. The first advancement was the explicit inclusion of positronic particles in order to describe the meta-stable states that positrons can form with usually polar molecules.[48, 49, 50, 38, 39] These systems of molecules and positrons are described considering a general set of $N_{f}$ fermions with charges and masses $\{q_{i},m_{i}\}_{i=1}^{N_{f}}$ and coordinate vectors $\bar{\textbf{x}}=\{\textbf{x}_{i}\}_{i=1}^{N_{f}}$, that include the three dimensional Cartesian and spin coordinates $\textbf{x}_{i}=(\textbf{r}_{i},\sigma_{i})$, and a set of $N_{n}$ nuclei with charges and masses $\{Z_{a},M_{a}\}_{a=1}^{N_{n}}$ and coordinate vectors $\bar{\textbf{R}}=\{\textbf{R}_{a}\}_{a=1}^{N_{n}}$.
@@ II.2 The Hamiltonian of Quantum Drude Oscillators
$\hat{h}_{i}^{O}=-\frac{1}{2\mu_{i}}\nabla^{2}_{\mathbf{r}_{i}^{d}}+v_{i}\left(r^{cd}_{i}\right)-\sum_{j(\neq i)=1}^{N_{O}}\frac{\rho_{i}\rho_{j}}{\left|\mathbf{r}_{i}^{d}-\mathbf{R}_{j}^{O}\right|},$ (4)
@@ II.3 Fermionic coupling to Quantum Drude Oscillators
An electronic system containing $N_{n}$ nuclei and $N_{f}$ fermions can be embedded into a bath of $N_{O}$ QDOs gives us the total Hamiltonian of the form
@@ II.4 External point charges as classical embedding
To simulate the effects on an atomic and positronic subsystem of an environment of molecules with an intrinsic dipole in the El-QDO embedding method[40, 42] together with the QDOs, it is possible to introduce fractional point charges[59] that act as a classical static Force-Field with a similar formalism to the one used in classical Force-Fields such as GAFF[68] and CHARMM[69].
Molecular Dynamics (MD) is not implemented in the code yet, and this will be one of the possible future developments. For now, a given set of point charges can be assigned to a specific molecule to describe intrinsic dipoles, like in the QDO model of water by Martyna and coworkers[59] first used in embedding procedures in ref. 40, 42. The interaction potential between the point charges and the systems of fermions and nuclei is written as
@@ II.5 Screening of the interactions
Both interaction potentials in eqs. 6 and 7 present divergences when a fermionic particle approaches the particles in the QDOs (centers and drudons) or the point charges present in the embedding potential, or when the drudons approach the molecular nuclei. In order to cure such divergences the total wavefunction of the system must include the description of all the cusps’ conditions as described in Section III.3.1. An alternative, and more efficient approach, that we use to screen the short-range interaction between the fermions, the two QDO particles and the point charges is that of introducing a damping function to remove the divergence of the Coulomb interactions. In ref. 42 the study of different screen functions has been presented. From this study it was evident that the error function damping
@@
where $\sigma_{ij}$ is specified for all pairs of particles using the combination rule $\sigma_{ij}=\sqrt{\sigma_{i}^{2}+\sigma_{j}^{2}}$, already used in previous works 59, 40, 42 was the most efficient to guarantee stability and convergence.
@@ II.6 Atomic pseudopotentials
QMeCha supports also the use of pseudopotentials to substitute the core electrons of atoms, reducing the number of degrees of freedom, and the electronic energy scaling involved in the calculations. The effective core potentials (ECPs) used, are those written in the standard semi-local form[70, *tra+14jctc, *tra+15jcp, *kro+16prb, *tra+17jcp, *tra+05jcp2, *ben+17jcp, *ben+18jcp, *ann+18jcp, *bur+07jcp, *bur+08jcp]
@@ III Trial wavefunctions
$\Psi_{T}(\bar{\textbf{x}},\bar{\textbf{r}}^{d})=\Psi_{f}(\bar{\textbf{x}})\Psi_{d}(\bar{\textbf{r}}^{d})\Psi_{f,d}(\bar{\textbf{r}},\bar{\textbf{r}}^{d}),$ (16)
which are respectively the wavefunction of the fermionic particles (electrons and positrons) $\Psi_{f}(\bar{\textbf{x}})$, the wavefunction of the purely drudonic particle, when present, $\Psi_{d}(\bar{\textbf{r}}^{d})$ and the coupling function between the fermionic system and the QDO environment, ie $\Psi_{f,d}(\bar{\textbf{r}},\bar{\textbf{r}}^{d})$[40]. A schematic representation of the trial wavefunction present in the code is shown in Fig. 2.
@@
The most general wavefunction $\Psi_{f}(\bar{\textbf{x}})$ for many electrons and many positrons must explicitly describe the many-body correlation effects between the $4N_{e}$ electronic Cartesian and spin coordinates $\bar{\textbf{x}}^{e}$ and the $4N_{p}$ positronic $\bar{\textbf{x}}^{p}$ ones in the field of the nuclei $\bar{\textbf{R}}$. Yet, this explicitly correlated form is in practice difficult to treat, and various approximations have been introduced in the literature, such as that proposed by Bressanini et al. in ref. 82, which is built as the product (or a linear combination of symmetrized products) of two-particle functions, including all particle pairs’ correlations (electron-electron, electron-positron, nucleus-electron, and nucleus-positron) correctly symmetrized (or anti-symmetrized) to take into account the distinguishability (or indistinguishability) of the particles with opposite (or identical) charge and spin.
@@
$\Psi_{f}(\bar{\textbf{x}})=\psi_{e}(\bar{\textbf{x}}^{e})\psi_{p}(\bar{\textbf{x}}^{p})\mathcal{J}(\bar{\textbf{x}})$ (17)
each depending only on the electronic $\psi_{e}(\bar{\textbf{x}}^{e})$ or positronic $\psi_{p}(\bar{\textbf{x}}^{p})$ degrees of freedom[83, 84, 85, 86, 87]. In this case, the correlation effects between electrons and positrons are only recovered through the Jastrow factor $J(\bar{\textbf{x}})$ (see Section III.3) that includes the description of all the nuclear and fermionic cusp conditions, together with various few-body terms usually limited to three- or four-body correlations, that are fundamental in describing the dynamical correlation between particles.
In the sections that follow we will describe the types of wavefunctions included in QMeCha that can be used with basis sets of atomic orbitals to describe $\psi_{e}(\bar{\textbf{x}}^{e})$ and $\psi_{p}(\bar{\textbf{x}}^{p})$. We will start with the most general one, ie the Pfaffian, to the most common Slater determinant, discussing the relationship between them[88, 89, 90, 91, 92].
@@ III.1.1 The Pfaffian
where the indexes $q$ and $p$ run on the full set of basis set orbitals $Q$, and $\left|1,1\right\rangle$ and $\left|1,-1\right\rangle$ correspond respectively to the triplet spin states $\left|\frac{1}{2},\frac{1}{2}\right\rangle\left|\frac{1}{2},\frac{1}{2}\right\rangle$ and $\left|\frac{1}{2},-\frac{1}{2}\right\rangle\left|\frac{1}{2},-\frac{1}{2}\right\rangle$. In order to fix the spin symmetry, the spatial part of the geminal functions must be antisymmetric with respect to the exchange of the electronic coordinates, thus the conditions that $\zeta^{\uparrow}_{qp}=-\zeta^{\uparrow}_{pq}$ and $\zeta^{\downarrow}_{qp}=-\zeta^{\downarrow}_{pq}$ must always hold. Moreover, since two electrons with parallel spin cannot occupy the same orbital, it must be that $\zeta^{\uparrow}_{qq}=0$ and $\zeta^{\uparrow}_{qq}=0$ $\forall\ q\in Q$. Because of these two conditions, it is clear that the set of coefficients $\zeta^{\uparrow}_{qp}$ and $\zeta^{\downarrow}_{qp}$ form two skew-symmetric matrices $\mathcal{Z}^{\uparrow}$ and $\mathcal{Z}^{\downarrow}$.
@@
where the coefficients $q$ and $p$, as for the triplet geminals in eqs. 21 and 22, run over the full basis set $Q$. In order to fix the spin symmetry, the spatial part of the geminal functions must be strictly symmetric, and thus the coefficients have to satisfy the condition $\lambda_{qp}=\lambda_{pq}$ and can be regrouped in a symmetric matrix $\Lambda$.
In the case in which the number of electrons (or positrons) in the system is odd, the Pfaffian wavefunction can be generalized following the procedure proposed by Sorella and coworkers in refs. 91, 92 or by simply adding a column (and a row) of elements, representing the electrons occupying an unpaired molecular orbital[88, 89]. In QMeCha we follow this last scheme, for which

### v2

@@ Abstract
We present the first open access version of the QMeCha (Quantum MeCha /’m”ke/) code, a quantum Monte Carlo (QMC) package developed to study many-body interactions between different types of quantum particles, with a modular and easy-to-expand structure. QMeCha is now available under CC BY-NC-ND license through the repository github.com/QMeCha. The present code has been built to solve the Hamiltonian of a system that can include nuclei and fermions of different mass and charge, e.g. electrons and positrons, embedded in an environment of classical charges and quantum Drude oscillators. To approximate the ground state of this many-particle operator, the code features different wavefunctions. For the fermionic particles, beyond the traditional Slater determinant, QMeCha also includes Geminal functions such as the Pfaffian, and presents different types of explicit correlation terms in the Jastrow factors. The classical point charges and quantum Drude oscillators, described through different variational ansätze, are used to model a molecular environment capable of explicitly describing dispersion, polarization, and electrostatic effects experienced by the nuclear and fermionic subsystem. To integrate these wavefunctions, efficient variational Monte Carlo and diffusion Monte Carlo protocols have been developed, together with a robust wavefunction optimization procedure that features correlated sampling.
@@ I Introduction
The most successful of these computational approaches fall into three main categories. These include wavefunction-based methods, such as Hartree-Fock (HF)[3, 4, 5, 6], Configuration Interaction[14, 15, 16] (CI) and Coupled Cluster[17, 18, 19, 20] (CC) multi-determinantal expansions, complete active-space (CAS) theories[21]; the long list of perturbative approaches like Møller-Plesset (MP) perturbation theory[22] and symmetry-adapted perturbation theory (SAPT)[23]; and the density-based methods, such as Density Functional Theory (DFT)[24, 25].
@@
The long list of modern QMC methods can be divided into two main categories. The first category includes the methods that are based on the stochastic sampling of the system’s degrees of freedom, such as the particles’ coordinates, and includes the most famous variational Monte Carlo (VMC)[26, 27], diffusion Monte Carlo (DMC)[28] and Lattice-Regularized diffusion Monte Carlo (LRMDC)[31] methods. The second category, on the other hand, contains all the QMC methods based on the sampling of the configurational space, such as Auxiliary-Field QMC[32], Full Configuration Interaction QMC (FCIQMC)[33, 34], and many others[28, 35].
Within this vast ensemble of QMC methods[36] many excellent open and restricted access QMC codes have been developed, such as TurboRVB[37], QMCPACK[38, 39], CASINO[40], QWALK[41], CHAMP[42], PyQMC[43], QMC=Chem[44] and HANDE-QMC[45].
In this work, we present the first version of the QMeCha (Quantum MeCha /’m”ke/)  [^1] The term mecha /’m”ke/ is the abbreviation of mechanical, which in Japanese manga and anime culture refers to all the piloted mechanical devices (robots, cars, airships, computers, and others) introduced in the science fiction genre. QMC package, initiated in 2017 as a flexible tool to study many-particle Hamiltonians with different wavefunctions’ approximations. In the years that followed, the code progressively grew to describe interactions between fermions of different charge and mass, such as positrons, muons and anti-muons, interacting with molecules[47, 48], and to include a QMC embedding procedure in which a molecular system of nuclei and fermions can be immersed in an environment approximated by point charges and quantum Drude oscillators (QDOs)[49, 50, 51]. The full Hamiltonian that comprises the two types of quantum particles, ie fermions and drudons, is integrated over one correlated wavefunction that is constructed as the product of various fermionic terms, such as the Slater determinant, and Jastrow factors. Within the code, the observables are computed through an efficient implementation of the VMC method, and through an improved size-consistent DMC algorithm, that reduces finite time-step errors when computing energy differences[36, 52, 53].
@@
Finally, QMeCha has been used to tackle three main sets of applications reported in Section VI. The first application focuses on the description of van der Waals (vdW) interactions in macromolecules[54], where the DMC calculations are used as state-of-the-art references together with the LNO-CCSD(T)[55, 56, 57, 58] method, to construct a dataset for the description of pocket-ligand interactions in proteins. The second application is related to positronic chemistry, an emerging field dedicated to the investigation of the still vastly unexplored bound metastable states of positrons with atomic and molecular systems, before electron-positron pair annihilation[59, 60, 61]. It has been shown that QMC methods are a promising tool for studying electron-positron bound states, because they can explicitly describe the electrostatic correlation effects between electron–positron pairs through wavefunctions including the two particle distances[62, 63, 64, 47, 65, 66, 67], unlike many quantum-chemical post-HF methods (like MP, CC, and CI) that have to rely on extremely large single-particle basis sets. The third set of applications is that of the El-QDO embedding method[49, 50, 51] first introduced in ref. 49 to efficiently capture the quantum correlation effects between an environment modelled through QDOs and point charges[68, 69, 70] and a molecular subsystem of interest. In this application we have focused on the evaluation of solvation energies, excitation energies and bond energy variations of a set of dimers in water clusters.[49, 51]
@@ II.1 Fermionic Hamiltonian
QMeCha was first implemented to study non-periodic molecular systems. The first advancement was the explicit inclusion of positronic particles in order to describe the meta-stable states that positrons can form with usually polar molecules.[59, 60, 61, 47, 48] These systems of molecules and positrons are described considering a general set of $N_{f}$ fermions with charges and masses $\{q_{i},m_{i}\}_{i=1}^{N_{f}}$ and coordinate vectors $\bar{\textbf{r}}^{f}=\{\textbf{r}_{i}\}_{i=1}^{N_{f}}$, that include the three dimensional Cartesian coordinates, and a set of $N_{n}$ nuclei with charges and masses $\{Z_{a},M_{a}\}_{a=1}^{N_{n}}$ and coordinate vectors $\bar{\textbf{R}}=\{\textbf{R}_{a}\}_{a=1}^{N_{n}}$.
@@ II.2 The Hamiltonian of Quantum Drude Oscillators
$\hat{h}_{i}^{O}=-\frac{1}{2\mu_{i}}\nabla^{2}_{\mathbf{r}_{i}^{d}}+v_{i}\left(r^{Od}_{i}\right)-\sum_{j(\neq i)=1}^{N_{O}}\frac{\rho_{i}\rho_{j}}{\left|\mathbf{r}_{i}^{d}-\mathbf{R}_{j}^{O}\right|},$ (4)
@@ II.3 Fermionic coupling to Quantum Drude Oscillators
An electronic system containing $N_{n}$ nuclei and $N_{f}$ fermions can be embedded into a bath of $N_{O}$ QDOs giving us the total Hamiltonian of the form
@@ II.4 External point charges as classical embedding
To simulate the effects on an atomic and positronic subsystem of an environment of molecules with an intrinsic dipole, such as the water clusters in Fig. 1, in the El-QDO embedding method[49, 51] together with the QDOs, it is possible to introduce fractional point charges[70] that act as a classical static Force-Field with a similar formalism to the one used in classical Force-Fields such as GAFF[79] and CHARMM[80].
Molecular Dynamics (MD) is not implemented in the code yet, and this will be one of the possible future developments. For now, a given set of point charges can be assigned to a specific molecule to describe intrinsic dipoles, like in the QDO model of water by Martyna and coworkers[70] first used in embedding procedures in ref. 49, 51.
Assuming that each pseudo-molecule in the embedding environment has associated only one QDO and many point charges, and that $n^{(i)}_{C}$ are the number of point charges associated to the $i$th pseudo-molecule corresponding to the $i$th QDO, then the full set of interactions within the embedding potential is written as:
$\text{V}_{O,C}=\sum_{j>i=1}^{N_{O}}\left[\sum_{a=1}^{n^{(i)}_{C}}\sum_{b=1}^{n^{(j)}_{C}}\frac{Q_{a}Q_{b}}{\left|\textbf{R}^{C}_{a}-\textbf{R}_{b}^{C}\right|}+\right.\\ \left.+\sum_{a=1}^{n^{(i)}_{C}}\left(\frac{Q_{a}\rho_{j}}{\left|\textbf{R}^{O}_{j}-\textbf{R}_{a}^{C}\right|}-\frac{Q_{a}\rho_{j}}{\left|\textbf{r}^{d}_{j}-\textbf{R}_{a}^{C}\right|}\right)\right],$ (7)
where the first terms represent the interactions between point charges of different pseudo-molecules, while the second and third terms represent the interactions between the point charges of one pseudo-molecule with the QDO charges of the other ones in the embedding environment.
The interaction potential between the point charges and the systems of fermions and nuclei is written as
@@ II.5 Screening of the interactions
Both interaction potentials in eqs. 6, 7, and 8 present divergences when a fermionic particle approaches the particles in the QDOs (centers and drudons) or the point charges present in the embedding potential, or when the drudons approach the molecular nuclei, and the point charges on the other QDOs. In order to cure such divergences the total wavefunction of the system must include the description of all the cusps’ conditions as described in Section III.3.1. Yet, the correct description of the cusps, is not able to properly account for the correct behaviour of the short range interactions between the various particles. Thus, an alternative, and more efficient approach, that we have used for now only to screen the short-range interaction between the fermions, the two QDO particles and the point charges is that of introducing a damping function to completely remove the divergence of the Coulomb interactions. In ref. 51 the study of different screening functions and the effect on the short-range interaction regime, has been presented. From this study, it was evident that the error function damping, used in previous works[70, 49, 51],
@@
where $q_{i}$$q_{j}$, represent two charges of different particles (nuclei, drudons, fermions) and $\sigma_{ij}=\sqrt{\sigma_{i}^{2}+\sigma_{j}^{2}}$ is the cut-off parameter determined by the single cut-offs $\sigma_{i}$ and $\sigma_{j}$ of the interacting charges, was the most efficient to guarantee stability and convergence.
@@ II.6 Atomic pseudopotentials
QMeCha supports also the use of pseudopotentials to substitute the core electrons of atoms, reducing the number of degrees of freedom, and the electronic energy scaling involved in the calculations. The effective core potentials (ECPs) used, are those written in the standard semi-local form[81, *tra+14jctc, *tra+15jcp, *kro+16prb, *tra+17jcp, *tra+05jcp2, *ben+17jcp, *ben+18jcp, *ann+18jcp, *bur+07jcp, *bur+08jcp]
@@ III Trial wavefunctions
$\Psi_{T}(\bar{\textbf{r}}^{f},\bar{\textbf{r}}^{d})=\Psi_{f}(\bar{\textbf{r}}^{f})\Psi_{d}(\bar{\textbf{r}}^{d})\Psi_{f,d}(\bar{\textbf{r}}^{f},\bar{\textbf{r}}^{d}),$ (17)
which are respectively the wavefunction of the fermionic particles (electrons and positrons) $\Psi_{f}(\bar{\textbf{r}}^{f})$, the wavefunction of the purely drudonic particle $\Psi_{d}(\bar{\textbf{r}}^{d})$, when present, and the coupling function between the fermionic system and the QDO environment, i.e. $\Psi_{f,d}(\bar{\textbf{r}}^{f},\bar{\textbf{r}}^{d})$[49]. A schematic representation of the trial wavefunction present in the code is shown in Fig. 2.
Within the code, no explicit dependency on the spin is present, and we omit to include those degrees of freedom. The vector $\bar{\textbf{r}}_{f}$, thus, represents only the $3N_{f}$ Cartesian coordinates of the electronic and positronic particles, while the $\bar{\textbf{r}}^{d}$ represents those of the $3N_{d}$ drudons. Therefore, the spin symmetry of the total wavefunction is only taken into account by imposing the correct symmetry of the spatial part of the wavefunction.
@@ III.1 Wavefunctions for electrons and positrons
The most general wavefunction $\Psi_{f}(\bar{\textbf{r}})$ for many electrons and many positrons must explicitly describe the many-body correlation effects between the $3N_{e}$ electronic Cartesian coordinates $\bar{\textbf{r}}^{e}$ and the $3N_{p}$ positronic $\bar{\textbf{r}}^{p}$ ones in the field of the nuclei $\bar{\textbf{R}}$. Yet, this explicitly correlated form is in practice difficult to treat, and various approximations have been introduced in the literature, such as that proposed by Bressanini et al. in ref. 93, which is built as the product (or a linear combination of symmetrized products) of two-particle functions, including all particle pairs’ correlations (electron-electron, electron-positron, nucleus-electron, and nucleus-positron) correctly symmetrized (or anti-symmetrized) to take into account the distinguishability (or indistinguishability) of the particles with opposite (or identical) charge and spin.
@@
$\Psi_{f}(\bar{\textbf{r}})=\psi_{e}(\bar{\textbf{r}}^{e})\psi_{p}(\bar{\textbf{r}}^{p})\mathcal{J}(\bar{\textbf{r}})$ (18)
each depending only on the electronic $\psi_{e}(\bar{\textbf{r}}^{e})$ or positronic $\psi_{p}(\bar{\textbf{r}}^{p})$ degrees of freedom[94, 95, 96, 97, 98]. In this case, the correlation effects between electrons and positrons are only recovered through the Jastrow factor $J(\bar{\textbf{r}})$ (see Section III.3) that includes the description of all the nuclear and fermionic cusp conditions, together with various few-body terms usually limited to three- or four-body correlations, that are fundamental in describing the dynamical correlation between particles.
In the sections that follow we will describe the types of wavefunctions included in QMeCha that can be used with basis sets of atomic orbitals to describe $\psi_{e}(\bar{\textbf{r}}^{e})$ and $\psi_{p}(\bar{\textbf{r}}^{p})$. We will start with the most general one, ie the Pfaffian, to the most common Slater determinant, discussing the relationship between them[99, 100, 101, 102, 103].
@@ III.1.1 The Pfaffian
where the indexes $q$ and $p$ run on the full set of basis set orbitals $Q$, and $\left|1,1\right\rangle$ and $\left|1,-1\right\rangle$ correspond respectively to the triplet spin states $\left|\frac{1}{2},\frac{1}{2}\right\rangle\left|\frac{1}{2},\frac{1}{2}\right\rangle$ and $\left|\frac{1}{2},-\frac{1}{2}\right\rangle\left|\frac{1}{2},-\frac{1}{2}\right\rangle$. In order to fix the spin symmetry, the spatial part of the geminal functions must be antisymmetric with respect to the exchange of the electronic coordinates, thus the conditions that $\zeta^{\uparrow}_{qp}=-\zeta^{\uparrow}_{pq}$ and $\zeta^{\downarrow}_{qp}=-\zeta^{\downarrow}_{pq}$ must always hold. Moreover, since two electrons with parallel spin cannot occupy the same orbital, it must be that $\zeta^{\uparrow}_{qq}=0$ and $\zeta^{\downarrow}_{qq}=0$ $\forall\ q\in Q$. Because of these two conditions, it is clear that the set of coefficients $\zeta^{\uparrow}_{qp}$ and $\zeta^{\downarrow}_{qp}$ form two skew-symmetric matrices $\mathcal{Z}^{\uparrow}$ and $\mathcal{Z}^{\downarrow}$.
@@
where the coefficients $q$ and $p$, as for the triplet geminals in eqs. 22 and 23, run over the full basis set $Q$. In order to fix the spin symmetry, the spatial part of the geminal functions must be strictly symmetric, and thus the coefficients have to satisfy the condition $\lambda_{qp}=\lambda_{pq}$ and can be regrouped in a symmetric matrix $\Lambda$.
In the case in which the number of electrons (or positrons) in the system is odd, the Pfaffian wavefunction can be generalized following the procedure proposed by Sorella and coworkers in refs. 102, 103 or by simply adding a column (and a row) of elements, representing the electrons occupying an unpaired molecular orbital[99, 100]. In QMeCha we follow this last scheme, for which